# Probabilistic and Differentiable Wireless Simulation with Geometric Transformers

### v1

@@ 3.3 Predictive modelling
Inverse problems. This differentiability makes such a surrogate model well-suited to solve inverse problems. For instance, we can use it for receiver localization. Given a 3D environment $F$, transmitters $\{t_{i}\}$, and corresponding signals $\{h_{i}\}$, we can find the most likely receiver position and orientation as $\hat{r}=\operatorname*{arg\,min}_{r}\sum_{i}\lVert h_{\theta}(F,t_{i},r)-h\rVert^{2}$. The minimization can be performed numerically through gradient descent, thanks to the differentiability of the Wi-GATr surrogate.
3.4 Probabilistic modelling
@@
While a predictive model of the signal can serve as a powerful neural simulator, it has two shortcomings. Solving an inverse problem through gradient descent requires a sizable computational cost for every problem instance. Moreover, predictive models are deterministic and do not allow us to model stochastic forward processes or express the inherent uncertainty in inverse problems.
Equivariant diffusion model. To overcome this, we draw inspiration from the inverse problem solving capabilities of diffusion models using guidance [13]. In this case, we formulate the learning problem as a generative modelling task of the joint distribution $p_{\theta}(F,t,r,h)$ between 3D environment mesh $F$, transmitter $t$, receiver $r$, and channel $h$, for a single transmitter-receiver pair. Concretely, we follow the DDPM framework and use a Wi-GATr model as score estimator (denoising network). By using an invariant base density and an equivariant denoising network, we define an invariant generative model. See Appendix B for a detailed description of our diffusion model and the discussion of some subtleties in equivariant generative modelling.
Unifying forward prediction and inverse problems as conditional sampling. A diffusion model trained to learn the joint density $p_{\theta}(F,t,r,h)$ does not only allow us to generate unconditional samples of wireless scenes, but also lets us sample from various conditionals: given a partial wireless scene, we can fill in the remaining details, in analogy to how diffusion models for images allow for inpainting. To achieve this, we use the conditional sampling algorithm proposed by Sohl-Dickstein et al. [48]: at each step of the sampling loop, we fix the conditioning variables to their known values before feeding them into the denoising network.
This algorithm lets us solve signal prediction (sampling from $p_{\theta}(h|F,t,r)$), receiver localization (from $p_{\theta}(r|F,t,h)$), geometry reconstruction (from $p_{\theta}(F_{u}|F_{k},t,r,h)$), or any other inference task in wireless scenes. We thus unify “forward” and “inverse” modelling in a single algorithm. Each approach is probabilistic, enabling us to model uncertainties. This is important for inverse problems, where measurements often underspecify the solutions.
In principle, the unconditional diffusion objective should suffice to enable test-time conditional sampling. In practice, we find that we can improve the conditional sampling performance with two modifications. First, we combine training on the unconditional diffusion objective with conditional diffusion objectives. For the latter, we randomly select tokens to condition on and evaluate the diffusion loss only on the remaining tokens. Second, we provide the conditioning mask as an additional input to the denoising model. See Appendix B for details.
@@
4 New datasets
While several datasets of wireless simulations and measurements exist [41, 57, 3, 4], they either do not include geometric information, are not diverse, are at a small scale, or the signal predictions are not realistic. To facilitate the development of machine learning methods with a focus on geometry, we generate two new datasets of simulated wireless scenes. [^2] We are preparing the publication of the datasets. Both feature indoor scenes and channel information generated with a state-of-the-art ray-tracing simulator [1] at a frequency of 3.5 GHz. They provide detailed characteristics for each path between Tx and Rx, such as gain, delay, angle of departure and arrival at Tx/Rx, and the electric field at the receiver itself, which allows users to compute various quantities of interest themselves. See Appendix C for more details.
Wi3R dataset. Our first dataset focuses on simplicity: each of 5000 floor plans has the same size and number of rooms, and all walls have the same material across layouts. They differ only in their layouts, which we take from Wi3Rooms [41], Tx positions, and Rx positions. In Appendix C we define training, validation, and test splits as well as an out-of-distribution set to test the robustness of different models.
WiPTR dataset. Next, we generate a more varied, realistic dataset based on the floor layouts in the ProcTHOR-10k dataset for embodied AI research [18]. We extract the 3D mesh information including walls, windows, doors, and door frames and assign 6 different dielectric materials for different groups of objects. Our dataset consists of 12k different floor layouts, split into training, test, validation, and OOD sets as described in Appendix C. Not only does WiPTR stand out among wireless datasets in terms of its level of detail and scale, but because it is based on ProcTHOR-10k, it is also suited for the integration with embodied AI research.
5 Experiments
5.1 Predictive modelling
@@
Table 2: Signal prediction results. We show the mean absolute error on the received power in dBm (lower is better, best in bold). Top: In-distribution performance. Middle: Generalization under symmetry transformations. Bottom: Generalization to out-of-distribution settings. In almost all settings, Wi-GATr is the highest-fidelity surrogate model.
We focus on the prediction of the time-averaged non-coherent received power $h=\sum_{p}|a_{p}|^{2}$, disregarding delay or directional information that may be available in real measurements. We train predictive surrogates $h_{\theta}(F,t,r)$ that predict the power as a function of the Tx position and orientation $t$, Rx position and orientation $r$, and 3D environment mesh $F$, on both the Wi3R and WiPTR datasets. All models are trained with reciprocity augmentation, i. e., randomly flipping Tx and Rx labels during training. This improves data efficiency slightly, especially for the transformer baseline.
In addition to our Wi-GATr model, described in Sec. 3, we train several baselines. The first is a vanilla transformer [54], based on the same inputs and tokenization of the wireless scene, but without the geometric inductive biases. Next, we compare to the $\mathrm{E}(3)$-equivariant SEGNN [8], though we were only able to fit this model into memory for the Wi3R dataset. In addition, we train a PLViT model, a state-of-the-art neural surrogate for wireless scenes [24] that represent wireless scenes as an image centered around the Tx position. Finally, we attempt to compare Wi-GATr also to WiNeRT [41], a neural ray tracer. However, this architecture, which was developed to be trained on several measurements on the same floor plan, was not able to achieve useful predictions on our diverse datasets with their focus on generalization across floor plans. Our experiment setup and the baselines are described in detail in Appendix D.
@@
Signal prediction. In Fig. 2 we illustrate the prediction task on a WiPTR floor plan. We show signal predictions for the simulator as well as for surrogate models trained on only 100 floor plans. Despite this floor plan not being part of the training set, Wi-GATr is able to capture the propagation pattern well, while the transformer and ViT show memorization artifacts.
In Tbl. 2 we compare surrogate models trained on the full Wi3R and WiPTR datasets. Both when interpolating Rx positions on the training floor plans as well as when evaluating on new scenes unseen during training, Wi-GATr offers the highest-fidelity approximation of the simulator. Wi-GATr as well as the equivariant baselines are by construction robust to symmetry transformations, while the performance of a vanilla transformer degrades substantially. All methods but SEGNN struggle to generalize to an OOD setting on the Wi3R dataset. This is not surprising given that the training samples are so similar to each other. On the more diverse WiPTR dataset, Wi-GATr is almost perfectly robust under domain shift.
Data efficiency. Next, we study the data efficiency of the different surrogates in Fig. 4. Wi-GATr is more data-efficient than any other method with the exception of the $\mathrm{E}(3)$-equivariant SEGNN, which performs similarly well for a small number of training samples. This confirms that equivariance is a useful inductive bias when data is scarce. But Wi-GATr scales better than SEGNN to larger number of samples, showing that our architecture combines the small-data advantages of strong inductive biases with the large-data advantages of a transformer architecture.
Inference speed. One of the advantages of neural surrogates is their test-time speed. Both Wi-GATr and a transformer are over a factor of 20 faster than the ground-truth ray tracer (see Appendix D).
@@
Receiver localization. Next, we show how differentiable surrogates let us solve inverse problems, focusing on the problem of receiver localization. We infer the Rx position with the predictive surrogate models by optimizing through the neural surrogate of the simulator as discussed in Sec. 3.3. The performance of our surrogate models is shown in Fig. 4 and Appendix D. [^3] Neither the SEGNN nor PLViT baselines are fully differentiable with respect to object positions when using the official implementations from Refs. [7, 24]. We were therefore not able to accurately infer the transmitter positions with these architectures. The two neural surrogates achieve a similar performance when only one or two transmitters are available, a setting in which the receiver position is highly ambiguous. With more measurements, Wi-GATr lets us localize the transmitter more precisely.
5.2 Probabilistic modelling
@@
Figure 5: Probabilistic modelling. We formulate various tasks as sampling from the unconditional or conditional densities of a single diffusion model. (a): Unconditional sampling of wireless scenes $p(F,t,r,h)$. (b): Receiver localization as conditional sampling from $p(r|F,t,h)$ for two different values of $h$ and $r$. (c): Geometry reconstruction as conditional sampling from $p(F_{u}|F_{k},t,r,h)$ for two different values of $h$, keeping $t$, $r$, $F_{k}$ fixed.
Next, we experiment with our probabilistic approach. We train diffusion models on the Wi3R dataset. In addition to a Wi-GATr model, we study a transformer baseline, as well as a transformer trained on the same data augmented with random rotations. Both models are trained with the DDPM pipeline with 1000 denoising steps and samples from with the DDIM solver [49]. Our setup is described in detail in Appendix D.
Signal prediction, receiver localization, and geometry reconstruction as conditional sampling. In our probabilistic approach, signal prediction, receiver localization, and geometry reconstruction are all instances of sampling from conditional densities: $h\sim p_{\theta}(h|F,t,r)$, $r\sim p_{\theta}(r|F,t,h)$, and $F_{u}\sim p_{\theta}(F_{u}|F_{k},t,r,h)$, respectively. We qualitatively show results for this approach in Figs. 1 and 5. All of these predictions are probabilistic, which allows our model to express uncertainty in ambiguous inference tasks. When inferring Rx positions from a single measurement, the model learns multimodal densities, as shown in the middle of Fig. 5. When reconstructing geometry, the model will sample diverse floor plans as long as they are consistent with the transmitted signal, see the right panel of Fig. 5. Additional results on signal and geometry prediction are given in Appendix D.2.
@@
We quantitatively evaluate these models through the variational lower bound on the log likelihood of test data under the model. To further analyze the effects of equivariance, we test the model both on canonicalized scenes, in which all walls are aligned with the $x$ and $y$ axis, and scenes that are arbitrarily rotated. The results in Tbl. 3 show that Wi-GATr outperforms the transformer baseline across all three tasks, even in the canonicalized setting or when the transformer is trained with data augmentation. The gains of Wi-GATr are particularly clear on the signal prediction and receiver localization problems.
6 Discussion
Wireless signal transmission through electromagnetic wave propagation is an inherently geometric and symmetric problem. We developed a class of neural surrogates grounded in geometric representations and strong inductive biases. They are based on the Wi-GATr backbone, consisting of a new tokenization scheme for wireless scenes together with an $\mathrm{E}(3)$-equivariant transformer architecture. The proposed backbone is applied in two ways to wireless tasks: first, as a differentiable “forward” prediction model that maps the features to the signals; second, as a probabilistic diffusion model that captures the joint and conditional distributions of features and channels. We employed these designs in experiments on received power prediction, receiver localization, and geometry reconstruction, where our Wi-GATr models enabled precise predictions, outperforming various baselines.
Our analysis is in many ways a first step. The range of materials in our datasets is limited and we only experimented with measurements of the non-coherent total received power, which is a stable signal, but offers less spatial information than measurements of the time delay or angular information. More importantly, we only considered idealized inference tasks. For instance, our receiver localization problem assumed perfect knowledge of the room geometry and materials.
Nevertheless, we hope that we were able to highlight the benefits of a geometric treatment of wave propagation modelling. Augmenting or replacing the image-based or general-purpose representations and architectures prevalent in wireless modelling with geometric approaches has the potential of improving data efficiency, performance, and robustness.

### v2

@@ 3.3 Applications beyond forward modelling
Inverse problems solved by gradient descent. The differentiability of surrogate models make them well-suited to solve inverse problems. For instance, we can use them for receiver localization. Given a 3D environment ${\bm{F}}$, transmitters $\{{\bm{x}}^{\text{tx}}_{i}\}$, and corresponding signals $\{h_{i}\}$, we can find the most likely receiver position and orientation as $\hat{{\bm{x}}^{\text{rx}}}=\operatorname*{arg\,min}_{{\bm{x}}^{\text{rx}}}\sum_{i}\lVert h_{\theta}({\bm{F}},{\bm{x}}^{\text{tx}}_{i},{\bm{x}}^{\text{rx}})-h_{i}\rVert^{2}$. The minimization can be performed numerically through gradient descent, using the gradients $\frac{\partial h_{\theta}}{\partial{\bm{x}}^{\text{rx}}}$.
@@
Probabilistic inference with diffusion models. While a predictive model of the signal can serve as a versatile neural simulator, it has two shortcomings. First, solving an inverse problem through gradient descent requires a sizable computational cost for every problem instance. Second, predictive models are deterministic and do not allow us to model stochastic forward processes or express the inherent uncertainty in inverse problems. To overcome this, we draw inspiration from the inverse problem solving capabilities of diffusion models using guidance (Chung et al., 2022; Gloeckler et al., 2024). We hereby follow the DDPM framework and use a Wi-GATr model as score estimator (denoising network) to learn a joint distribution $p_{\theta}({\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$ between 3D environment mesh ${\bm{F}}$, transmitter ${\bm{x}}^{\text{tx}}$, receiver ${\bm{x}}^{\text{rx}}$, and channel $h$, for a single transmitter-receiver pair (see Fig. 1b).
4 New datasets with diverse scene geometry
To show the importance of geometry and symmetries in wireless signal propagation, we require a dataset that comprises complex signal interactions with diverse geometries. While several datasets of wireless simulations and measurements exist (Orekondy et al., 2022b; Zhang et al., 2023; Alkhateeb, 2019; Alkhateeb et al., 2023), they either do not include geometric information, lack scene diversity or the signal predictions are not realistic. Therefore, we generate two datasets that feature indoor scenes and channel information at a frequency of 3.5 GHz using a state-of-the-art ray-tracing simulator (Remcom, ). [^2] We are preparing the publication of the datasets. We focus on indoor scenes as transmission plays a stronger role than outdoors. The datasets provide detailed characteristics for each path between Tx and Rx, such as gain, delay, angle of departure and arrival at Tx/Rx, and the electric field at the receiver itself, which allows users to compute various quantities of interest themselves. See Appendix D for more details.
Wi3R dataset. Our first dataset focuses on diversity of room layouts and their impact on the signal. We take 5000 layouts from Wi3Rooms (Orekondy et al., 2022b) and randomly sample 3D Tx positions and Rx positions. In Appendix D we provide more details and define training, validation, and test splits as well as an out-of-distribution set to test the robustness of different models.
WiPTR dataset. To increase complexity, we generate a second, more varied, realistic dataset that includes material properties, varying number of faces, and varying layout dimensions. The floor layouts are based on the ProcTHOR-10k dataset for embodied AI research (Deitke et al., 2022). This dataset consists of 12k different floor layouts, split into training, test, validation, and OOD sets as described in Appendix D. WiPTR stands out among wireless datasets in terms of geometric diversity of scenes, which enables the geometric deep learning community to study symmetries in electromagnetic signal propagation. Additionally, since it is based on ProcTHOR-10k, it can provide wireless signals as a novel modality for embodied AI research.
5 Experiments
@@
In our experiments, we provide empirical evidence for the benefits of the inductive bias of our equivariant architecture, as well as the inference speed, differentiability, and expressiveness of fully-learnt neural simulation surrogates on simulated and real-world data. In addition, we demonstrate probabilistic inference through a diffusion model.
5.1 Signal strength prediction on simulated data
On the simulated data, we focus on the prediction of the time-averaged non-coherent received power $h=\sum_{p}|a_{p}|^{2}$, where $|a_{p}|^{2}$ is the received power of each independent ray tracing path $p$. We train surrogates $h_{\theta}({\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}})$ that predict the power as a function of the Tx position and orientation ${\bm{x}}^{\text{tx}}$, Rx position and orientation ${\bm{x}}^{\text{rx}}$, and 3D environment mesh ${\bm{F}}$, on both the Wi3R and WiPTR datasets. All models are trained with reciprocity augmentation, i. e., randomly flipping Tx and Rx labels during training. This improves data efficiency, especially for the transformer baseline.
Baselines. In addition to our Wi-GATr model, described in Sec. 3, we train several baselines. The first is a vanilla transformer (Vaswani et al., 2017), based on the same inputs and tokenization of the wireless scene, but without the geometric inductive biases. To ablate the effect of our tokenization scheme, we also test a transformer without it, which represents the scene as a simple sequence of 3D positions. Next, we compare to the $\mathrm{E}(3)$-equivariant SEGNN (Brandstetter et al., 2022b), though we were only able to fit this model into memory for the Wi3R dataset. As state-of-the-art method for image-based wireless channel modelling, we train a PLViT model (Hehn et al., 2023). We have also considered WiNeRT (Orekondy et al., 2022b) and Sionna RT (Hoydis et al., 2022) as baselines, but found them unsuited for the problem setup as explained in Appendix E. More details on our experiment setup and the baselines are also given in Appendix E.
In-distribution and out-of-distribution performance. In Fig. 2 we illustrate the prediction task on a WiPTR floor plan. We show signal predictions for the simulator as well as for surrogate models trained on only 100 floor plans. while the transformer and ViT show memorization artifacts, Wi-GATr is able to capture the propagation pattern, although this floor plan was not seen during training.
In Tbl. 1 we compare surrogate models trained on the full Wi3R and WiPTR datasets. Both when interpolating Rx positions on the training floor plans as well as when evaluating on new scenes unseen during training, Wi-GATr offers the highest-fidelity approximation of the simulator. Wi-GATr as well as the equivariant baselines are by construction robust to symmetry transformations, while the performance of a vanilla transformer degrades substantially. All methods but SEGNN struggle to generalize to an OOD setting on the Wi3R dataset. This is not surprising given that the training samples are so similar to each other. On the more diverse WiPTR dataset, Wi-GATr is almost perfectly robust under domain shift.
Data efficiency. Next, we study the data efficiency of the different surrogates in Fig. 4. Wi-GATr is more data-efficient than any other method with the exception of the $\mathrm{E}(3)$-equivariant SEGNN, which performs similarly well for a small number of training samples. This confirms that equivariance is a useful inductive bias when data is scarce. But Wi-GATr scales better than SEGNN to larger number of samples, showing that our architecture combines the small-data advantages of strong inductive biases with the large-data advantages of a transformer architecture. We find clear benefits of our wireless tokenizer when comparing the transformer baselines with and without the tokenizer.
Inference speed. One of the advantages of neural surrogates is their test-time speed. Both Wi-GATr and a transformer are over a factor of 20 faster than the ground-truth ray tracer (see Appendix E).
@@ 5.1 Signal strength prediction on simulated data
Table 1: Signal prediction results. We show the mean absolute error on the received power in dB (lower is better, best in bold). Top: In-distribution performance. Middle: Generalization under symmetry transformations. Bottom: Generalization to out-of-distribution settings. In almost all settings, Wi-GATr is the highest-fidelity surrogate model.
@@
5.2 Inverse receiver localization
@@
We show how differentiable surrogates let us solve inverse problems, focusing on the problem of receiver (Rx) localization. We infer the Rx position with the predictive surrogate models by optimizing through the neural surrogate of the simulator as discussed in Sec. 3.3. The performance of our surrogate models is shown in Fig. 4 and Appendix E. [^3] Neither the SEGNN nor PLViT baselines are fully differentiable with respect to object positions when using the official implementations from Refs. (Brandstetter et al., 2022a; Hehn et al., 2023). We were therefore not able to accurately infer the transmitter positions with these architectures. The two neural surrogates achieve a similar performance when only one or two transmitters are available, a setting in which the receiver position is ambiguous. With more measurements, Wi-GATr lets us localize the transmitter more precisely.
5.3 Invariant diffusion for inverse problems
@@
Figure 5: Probabilistic modelling. We formulate various tasks as sampling from the unconditional or conditional densities of a single diffusion model. (a): Unconditional sampling of wireless scenes $p({\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$. (b): Receiver localization as conditional sampling from $p({\bm{x}}^{\text{rx}}|{\bm{F}},{\bm{x}}^{\text{tx}},h)$ for two different values of $h$ and ${\bm{x}}^{\text{rx}}$. (c): Geometry reconstruction as conditional sampling from $p({\bm{F}}_{u}|{\bm{F}}_{k},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$ for two different values of $h$, keeping ${\bm{x}}^{\text{tx}}$, ${\bm{x}}^{\text{rx}}$, ${\bm{F}}_{k}$ fixed.
In the following, we consider the problem of signal prediction as well as the inverse problems of receiver localization and geometry reconstruction. All three are instances of sampling from conditional densities: $h\sim p_{\theta}(h|{\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}})$, ${\bm{x}}^{\text{rx}}\sim p_{\theta}({\bm{x}}^{\text{rx}}|{\bm{F}},{\bm{x}}^{\text{tx}},h)$ and ${\bm{F}}\sim p_{\theta}({\bm{F}}|{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$, respectively. We obtain the conditional densities by inpainting during the sampling from the joint densities (Sec. 3.3;Sohl-Dickstein et al. (2015)). To do so, we train a diffusion model using a Wi-GATr backbone with the DDPM pipeline and 1000 denoising steps. During training, we additionally apply conditional masking strategies to improve inpainting performance. See Appendix C for a detailed description of our diffusion model, the masking protocol, as well a discussion on equivariant generative modelling. For comparison, we study a transformer baseline and a transformer trained on the same data augmented with random rotations. More details on the model architectures can be found in Appendix E.
@@
We qualitatively show results for this approach in Figs. 1 and 5. All of these predictions are probabilistic, which allows our model to express uncertainty in ambiguous inference tasks. When inferring Rx positions from a single measurement, the model learns multimodal densities, as shown in the middle of Fig. 5. When reconstructing geometry, the model will sample diverse floor plans as long as they are consistent with the transmitted signal, see the right panel of Fig. 5. Additional results on signal and geometry prediction are given in Appendix E.2.
@@
We quantitatively evaluate these models through the variational lower bound on the log likelihood of test data under the model. To further analyze the effects of equivariance, we test the model both on canonicalized scenes, in which all walls are aligned with the $x$ and $y$ axis, and scenes that are arbitrarily rotated. The results in Tbl. 2 show that Wi-GATr outperforms the transformer baseline across all three tasks, even in the canonicalized setting or when the transformer is trained with data augmentation. The gains of Wi-GATr are particularly clear on the signal prediction and receiver localization problems.
5.4 Comparison to differentiable ray tracing on real-world data
The utility of wireless channel simulation is ideally evaluated on real data. Unfortunately, the scarcity of available measurement data limits evaluation to simple scenarios and prevents thorough testing of the geometric generalization capabilities. Yet, we aim to highlight limitations of hybrid ray tracing methods compared to our fully-learnt approach. To this end, we replicate the evaluation setup of Hoydis et al. (2023) on the DICHASUS dataset (Euchner et al., 2021).
Figure 6: Signal prediction on real measurements. As a function of the training data size, we show the mean absolute (logarithmic) error on the received power (left) and the relative absolute error on the delay spread (right).
The dataset consists of a single hallway with receiver arrays at each end and an occluded corridor in between. During collection, a robot-mounted transmitter was moving through the hallway, covering most of the area, with accurate location data coming from a tachymeter. We train our model to predict the average received power and the average delay spread for each receiver array, using the same data splits and evaluation scheme as in Hoydis et al. (2023). We report the mean absolute (logarithmic) error on the average received power and the relative absolute error on the average delay spread. For this purpose, we can consider each array as a single receiver without modifications to our Wi-GATr backbone. As baselines, we run their hybrid approaches named Learned Materials and Neural Materials.
In Figure 6, we show the performance in both prediction tasks with respect to the number of transmitter locations used during training. Each transmitter location corresponds to two data samples, one per receiver. The Learned Materials approach maintains consistent performance due to model bias, making it robust with limited data but less flexible with larger datasets. Neural Materials also benefits from model bias with little data but, thanks to its flexible material model, can leverage data to overcome other model and geometry limitations. Our fully-learnt approaches Wi-GATr and the Transformer outperform both hybrid models as the dataset size grows, while their drawback of higher variance in the small data regime can be stabilized through pretraining on WiPTR (indicated by “pt”). We find that using a Transformer with the Wi-GATr tokenization scheme slightly outperforms GATr in the small data regime. However, we attribute this to simplicity of the dataset, where symmetry transformations of 3D space are neither observed nor required.
6 Conclusion
We developed a class of neural surrogates grounded in geometric representations and strong inductive biases. They are based on the Wi-GATr backbone, consisting of a new tokenization scheme for wireless scenes together with an $\mathrm{E}(3)$-equivariant transformer architecture. In our experiments, we demonstrated the benefits of the inductive bias, inference speed, differentiability, and expressiveness of our fully-learnt equivariant neural simulation surrogate on simulated and real-world data. In addition, we showed how our backbone can be used in a diffusion model for probabilistic inference.
Limitations. Our paper presents a step towards fully-learnt wireless simulations, which has natural trade-offs when compared to model-based and hybrid ray tracers. Unlike model-based approaches (which require no training data), learnable approaches depend on reasonably-sized amounts of training data and hence involve data collection overhead. Furthermore, model-based approaches are highly configurable (e. g., scene parameterization, antenna patterns, controlling precision), whereas our approach is less configurable with many configuration choices learnt implicitly using training data. Finally, we emphasize that we are not proposing a drop-in replacement of a model-based ray tracer, but rather take a step towards the overarching goal of learning wireless simulations from data.
Acknowledgements.